OpenStudy (anonymous):

(a)n=[(n-1)/(n^2+1)], calculate the 6 values of limbs, also Lim(o-infiniti)(an)

OpenStudy (anonymous):

\[a_n=\frac{n-1}{n^2+1}\]?

OpenStudy (anonymous):

ja

OpenStudy (anonymous):

then \[\lim_{n\to \infty}a_n=0\]since the degree of the numerator is less than the degree of the denominator not sure what limbs are

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

LIMBS IS chain

OpenStudy (mathmate):

Does it mean to "calculate \( a_0\), \( a_1\), \( a_2\), ...\( a_5\), and \( Lim_0^\infty a_n \) " ?

OpenStudy (anonymous):

ja thats it

OpenStudy (mathmate):

For n=0, a0=(0-1)/(0+1)=-1 n=1, a1=(1-1)/(1+1)=0 .... until n=5, a5=(5-1)/(5^2+1)=4/26=2/13